# PDE-Transformer: A Continuous Dynamical Systems Approach to Sequence Modeling

### v2

@@ 2 Related Work
We situate the PDE-Transformer within three major research streams: (i) efficiency enhancements to discrete Transformers, (ii) continuous-time sequence modeling, and (iii) neural networks as PDE solvers. Together, these perspectives highlight both the progress and the remaining gaps that motivate our work.
2.1 Discrete Efficiency Enhancements
The quadratic complexity of the vanilla Transformer Vaswani et al. (2017) has motivated extensive efforts to improve scalability. One class of methods reduces attention cost by introducing fixed or learned sparsity, such as local windows combined with global tokens Beltagy et al. (2020); Zaheer et al. (2020) or strided/dilated patterns Child et al. (2019). While effective, these approaches impose pre-defined structures that may not align with data and risk bottlenecks in capturing long-range dependencies. A second class leverages low-rank or kernel approximations: Linformer Wang et al. (2020) projects Key and Value matrices into lower-dimensional spaces, while Performer Choromanski et al. (2021) approximates softmax attention with kernel features. These methods reduce complexity but rely on restrictive low-rank assumptions that can discard fine-grained information in the long-tail spectrum. A third family introduces recurrence and memory, such as Transformer-XL Dai et al. (2019), which caches hidden states across segments to extend context length. This alleviates quadratic cost but introduces challenges in compressing historical states and handling discontinuities at segment boundaries. Overall, these methods optimize the discrete computation graph—deciding “who attends to whom”—without altering the fundamental discrete paradigm (see Table 1).
Table 1: Taxonomy of efficient Transformer architectures. These methods improve scalability within a discrete framework but do not change the underlying paradigm.
Approach Key Examples Core Idea Limitation
Sparsity BigBird Zaheer et al. (2020), Longformer Beltagy et al. (2020) Restrict attention to sparse patterns Rigid information flow; may miss dependencies
Low-Rank / Kernel Linformer Wang et al. (2020), Performer Choromanski et al. (2021) Approximate the full attention matrix Low-rank assumption; loss of high-frequency info
Recurrence / Memory Transformer-XL Dai et al. (2019) Segment sequence and reuse past states State compression bottleneck; boundary effects
2.2 Continuous-Time Models
A more fundamental line of research interprets neural networks as discretizations of continuous-time systems. Neural ODEs Chen et al. (2018) recast residual networks as solutions of ODEs, treating depth as continuous and enabling benefits such as memory-efficient training Gholami et al. (2019) and irregular time-series modeling Li et al. (2020). However, these methods continuous-ize the depth dimension rather than the sequence.
State-Space Models (SSMs) directly target the sequence dimension. They model dynamics with a linear ODE:
$h^{\prime}(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t)+\mathbf{D}x(t),$ (1)
where $x(t)$ is the input, $h(t)$ the latent state, and $y(t)$ the output. The S4 model Gu et al. (2022) demonstrated that by structuring $\mathbf{A}$, one can discretize this system into an efficient convolutional form. Recent advances like Mamba Gu & Dao (2023) incorporate selection mechanisms and input-dependent parameters, surpassing linear time-invariant constraints and achieving strong performance. SSMs thus represent the most competitive continuous-time alternative to Transformers. However, they are governed by ODEs, which describe the temporal evolution of state vectors. Our PDE formulation instead models a field $u(x,t)$, introducing explicit spatial derivative terms (e.g., $\nabla^{2}u$) and enabling spatiotemporal coupling along the sequence dimension (Table 2).
Table 2: Comparison of continuous-time modeling paradigms. Unlike ODE-based models, our PDE approach introduces explicit spatiotemporal coupling.
Paradigm Continuous Dim. Equation Mechanism Limitation
Neural ODE Chen et al. (2018) Depth ODE Continuous layers Models depth, not sequence
SSM (Mamba) Gu & Dao (2023) Sequence ODE Temporal evolution + selection Lacks explicit spatial coupling
PDE-Transformer (Ours) Sequence PDE Spatiotemporal coupling Models direct interaction via $\nabla^{2}u$
2.3 Neural Networks as PDE Solvers
A parallel literature connects neural networks and PDEs but with inverse objectives. Physics-Informed Neural Networks (PINNs) Raissi et al. (2019) parameterize PDE solutions with neural networks, enforcing boundary conditions and PDE residuals during training. Neural Operators such as the Fourier Neural Operator (FNO) Li et al. (2021) extend this idea by learning resolution-invariant mappings from input functions to PDE solutions, effectively serving as universal solvers. In contrast, our work does not seek to solve externally-given PDEs. Instead, we adopt PDE principles—diffusion, waves, and reaction—as internal design guidelines for the information dynamics of sequence models.
2.4 Synthesis
In summary, discrete efficiency methods enhance scalability without altering the static computation graph; continuous-time models, particularly SSMs, provide strong alternatives but are ODE-based and lack explicit spatial coupling; and neural PDE solvers apply PDEs as external constraints rather than intrinsic dynamics. This landscape reveals an open frontier: applying PDEs as first-principles foundations for neural sequence architectures. Our PDE-Transformer advances this direction by formalizing spatiotemporal coupling as a core design principle for long-range modeling.
3 Theory and Method: Unified Variational Dynamical System and Adaptive Diffusion
To address the dual bottlenecks of quadratic complexity and the lack of local geometric modeling in Transformers, we reframe sequence modeling as a variational dynamical system from first principles. In this view, representations evolve under local diffusion, nonlinear reaction, and nonlocal coupling forces. This reveals a structural deficiency in standard Transformers—the missing diffusion term—which we remedy with a practical and efficient adaptive PDE diffusion layer. In this section, we derive the governing dynamics from an energy functional, map its components to Transformer modules, and instantiate the theory as a trainable layer, further proposing a principled integration strategy and discussing its multiscale spectral properties.
3.1 Unified Variational Principle
(a) Unified variational framework: energy functional decomposition and gradient-flow PDE.
(b) Architecture mapping: PDE components vs. standard Transformer, showing the missing diffusion term and our proposed PDE diffusion layer.
Figure 1: Theory framework of PDE-Transformer. Two stacked panels show (a) the unified variational formulation and corresponding PDE, and (b) the architectural mapping that highlights the missing diffusion component in standard Transformers.
We denote the continuous representation of the sequence as $u(x,t)$, where $x\in\Omega$ represents the position and $t$ the “evolution time” corresponding to network depth. The global energy functional is defined as
$E[u]$ $=\int_{\Omega}\Big(\tfrac{\alpha}{2}|\nabla u|^{2}+F(u)\Big)dx+\frac{\beta}{2}\iint_{\Omega\times\Omega}K(x,y)[u(x)-u(y)]^{2}dxdy.$ (2)
It consists of three parts: a local tension term $|\nabla u|^{2}$ that penalizes sharp variations to maintain local consistency; a nonlinear potential $F(u)$ encoding task objectives; and a long-range coupling $K(x,y)$ establishing content-dependent global interactions.
Theorem 3.1 (Unified Dynamical Equation).
The gradient flow of $E[u]$ is
$\frac{\partial u}{\partial t}=\alpha\Delta u-F^{\prime}(u)-\beta\mathcal{L}_{K}[u],\qquad\mathcal{L}_{K}[u](x)=\int K(x,y)[u(x)-u(y)]dy.$ (3)
Each term corresponds directly to a Transformer module. The standard Transformer realizes reaction and nonlocal coupling but ignores diffusion, limiting explicit geometric modeling and stability.
Table 3: Physical and functional correspondence of PDE-Transformer components.
Mathematical Term Physical Process Module Function Information Pattern
$\alpha\Delta u$ Diffusion PDE Diffusion Layer (Ours) Regularizes locally Geometric, dense
$F^{\prime}(u)$ Reaction Feedforward Network Pointwise transformation Local, independent
$\beta\mathcal{L}_{K}[u]$ Nonlocal coupling Self-Attention Aggregates globally Sparse, content-driven
3.2 Diffusion Dynamics and Stability
To align theory with implementation, we adopt Neumann (reflective) boundary conditions:
$(\mathbf{\Delta}_{N}X)_{i}=\begin{cases}X_{2}-X_{1},&i=1,\\ X_{i-1}-2X_{i}+X_{i+1},&2\leq i\leq L-1,\\ X_{L-1}-X_{L},&i=L.\end{cases}$ (4)
This preserves boundary information and is compatible with discrete cosine transform (DCT) analysis.
Theorem 3.2 (Spectrum).
The eigenvalues of $\mathbf{\Delta}_{N}$ are $\lambda_{k}=-4\sin^{2}(\pi k/(2L))$ with cosine eigenvectors.
Corollary 3.3 (CFL Condition).
For the explicit iteration $X^{(k+1)}=X^{(k)}+\alpha\mathbf{\Delta}_{N}X^{(k)}$, stability requires $\alpha<\tfrac{1}{2}$.
Theorem 3.4 (Lyapunov Monotonicity).
Under the CFL condition, the Dirichlet energy $V(X)=X^{\top}(-\mathbf{\Delta}_{N})X$ decreases monotonically, ensuring stability and improving the optimization landscape.
3.3 Information-Theoretic Integration
Consider a forward chain $X_{0}\to X_{1}\to\cdots\to X_{N}\to Y$. For any smoothing operator $\mathcal{S}$, define the information retention:
$\rho(d)=\mathbb{E}\left[\frac{I(\mathcal{S}(X^{(d)});Y)}{I(X^{(d)};Y)}\right],\qquad\rho(1)\geq\rho(2)\geq\cdots\geq\rho(N).$ (5)
This shows that earlier diffusion preserves more task-relevant information.
We further compare seven candidate insertion points using the value function $V_{i}=w_{I}I_{i}-w_{D}D_{i}-w_{C}C_{i}$, where $I_{i}$ denotes information retention, $D_{i}$ distortion, and $C_{i}$ computational cost.
Table 4: Comparison of candidate insertion points for diffusion layers.
Pos. Mechanism Property Info Dist. Cost Rank
1 After Embedding $\|X_{0}^{(s)}-X_{0}\|=O(\alpha)$ High Low Low 1
2 After MLP Local proximal step High Low Low 2
3 Between Layers Gradient flow improvement Mid-High Mid Mid 3
4 Before LN Non-commutativity Mid Mid Low 4
5 Inside Attention Breaks normalization Mid-Low Mid-High Mid 5
6 Between Heads Loss of diversity Low High Mid 6
7 After Attention Breaks sparsity Low High Low 7
Overall Ranking: After embedding >> After MLP >> Between layers >> Before LN >> Inside attention >> Between heads >> After attention.
3.4 Multi-Scale Spectral Complementarity
Figure 2: Frequency domain analysis of the multi-scale diffusion mechanism. (Left) The transfer function $H(\omega)$ for different diffusion scales. Single-scale diffusions (Fast, Medium, Slow) act as low-pass filters with different cutoff frequencies. (Right) The energy distribution across four frequency bands. The multi-scale approach (green, dashed line) achieves a more balanced energy distribution across the entire frequency spectrum compared to any single-scale method, enabling it to capture a richer set of signal components from both global trends (low frequency) and local details (high frequency).
For diffusion with step size $h$, the frequency response is
$H_{h}(\omega)=1-4\alpha\sin^{2}\!\Big(\tfrac{\omega h}{2}\Big).$ (6)
Small steps ($h=1$) preserve high-frequency details, while larger steps ($h=2,4$) emphasize mid-to-low frequencies. Mixing scales covers the entire spectrum; a practical initialization uses weights $1:0.6:0.3$. Diffusion is geometry-driven and dense, with effective radius proportional to $\sqrt{t}$, while attention is content-driven and sparse, connecting arbitrary positions. Their combination captures both local geometry and global dependency.
3.5 Adaptive PDE Diffusion Layer
The proposed module incorporates multi-scale diffusion ($[1,2,4]$), learnable coefficients per scale and channel, and CFL-enforced constraints $\alpha<0.5$ at runtime. Its complexity is $O(Ld)$ for a single scale or $O(KLd)$ for multiple scales, significantly lower than attention $O(L^{2}d)$. Theory and information analysis indicate that the best insertion point is after the embedding layer, before the first Transformer block, maximizing information retention and stability while improving optimization.
Summary. We unify sequence modeling into a reaction–diffusion–nonlocal coupling system, derive its integro-PDE gradient flow, and identify the missing diffusion term in Transformers. Spectral and stability analysis motivates an efficient adaptive PDE diffusion layer, with its optimal position after embedding. This principled module augments global coupling with explicit local modeling, laying the foundation for subsequent experiments. applications.

### v3

@@ 2 Related Work
Our work sits at the intersection of (i) long-sequence Transformers, (ii) locality-inducing inductive biases, and (iii) stability- and operator-motivated architectural analysis. We position our contribution along three axes: we do not change attention’s routing rule or complexity (unlike efficient-attention variants), we use PDE diffusion as a fixed-form, spectrally-budgeted local prior (not to be confused with generative diffusion models), and we provide a structural explanation for insertion-position sensitivity (often treated empirically in hybrid designs).
2.1 Long-Sequence Transformers
The quadratic complexity of self-attention has motivated extensive work on efficient long-sequence Transformers. Representative directions include sparse patterns (e.g., Longformer (Beltagy et al., 2020), BigBird (Zaheer et al., 2020), Sparse Transformers (Child et al., 2019), Reformer (Kitaev et al., 2020)), low-rank or kernelized approximations (e.g., Performer (Choromanski et al., 2020), Linear Transformers (Katharopoulos et al., 2020), Linformer (Wang et al., 2020), Nyströmformer (Xiong et al., 2021)), and memory-optimized exact attention (e.g., FlashAttention (Dao et al., 2022)). LRA (Tay et al., 2020b, a) further highlights that efficiency improvements do not automatically translate into robust long-context generalization.
Our work is orthogonal: we keep attention intact and instead study how to inject an explicit local geometric prior via a lightweight PDE diffusion operator. Because the PDE diffusion layer is linear-time, it can be combined with sparse/linear attention mechanisms as a complementary design component.
2.2 Diffusion, PDEs, and Neural Operators
PDE- and diffusion-inspired views have a long history in deep learning, including stability-aware architectures via dynamical-systems discretizations (Haber & Ruthotto, 2017; Ruthotto & Haber, 2018) and classical diffusion for denoising/smoothing (Weickert, 1998). More recently, neural operator learning aims to approximate families of PDE solution operators (Li et al., 2020), and generative diffusion models treat diffusion as a stochastic generative process (Ho et al., 2020).
Our use of PDE diffusion differs in both goal and modeling choice. We use a fixed-form discrete Laplacian with a strict spectral budget to inject a local smoothness prior with unconditional stability guarantees, rather than learning an operator approximator or modeling a generative diffusion process. The resulting module is a deterministic architectural regularizer that targets the sequence dimension and integrates naturally with Transformers.
2.3 Spectral Stability and Regularization in Deep Networks
Stability and spectral control have been recognized as central issues in deep learning, particularly for very deep or long-context models. Spectral normalization (Miyato et al., 2018) and Lipschitz-constrained architectures aim to control gradient explosion and improve robustness. Layer normalization (Ba et al., 2016) and residual connections (He et al., 2015) provide fundamental stability mechanisms in deep networks. In Transformers, prior work has analyzed the instability of attention maps and the amplification of high-frequency components across layers.
Our work contributes to this line by introducing a simple yet effective mechanism for spectral control along the sequence dimension. By leveraging the discrete Laplacian with a strict spectral budget, we guarantee non-expansiveness and monotone Dirichlet energy dissipation for the diffusion subsystem. Importantly, our analysis does not rely on global Lipschitz assumptions for the full Transformer; instead, it isolates what can be proven unconditionally and studies interactions with attention through controlled perturbation and non-commutativity.
2.4 Position Sensitivity and Hybrid Architectures
Hybrid architectures that combine local and global operations are increasingly common, including convolution–attention hybrids (Wu et al., 2019; Dai et al., 2020) and alternatives to attention such as structured state space models (Gu et al., 2021; Gu & Dao, 2023). In many such designs, the placement of local operators relative to global mixing is decided empirically.
Our contribution is a principled account of this placement sensitivity. By framing diffusion and attention as generally non-commuting operators, we explain why pre-routing diffusion can act as beneficial manifold regularization while post-routing diffusion can introduce a destructive double-mixing effect. This operator-theoretic lens is not specific to diffusion and can serve as a transferable template for reasoning about other local–global compositions.
3 Operator Framework and Diffusion Layer Design
3.1 Scope of Theoretical Guarantees
This section introduces an operator-theoretic framework centered on a PDE diffusion layer and clarifies the precise scope of its theoretical guarantees. Our analysis is intentionally structured into three conceptual layers, each corresponding to a different level of rigor and assumption.
At the first level, we establish intrinsic stability properties of the diffusion operator itself. These include spectral non-expansiveness and monotone energy dissipation, which hold unconditionally and independently of all other Transformer components. These results are provable without assumptions on attention, normalization, or nonlinear activations.
At the second level, we study how diffusion interacts compositionally with backbone modules such as LayerNorm, self-attention, and MLP blocks. This analysis relies on local Lipschitz assumptions and yields perturbation propagation bounds that explain why the effect of diffusion depends strongly on its insertion position within the network.
At the third level, we provide a diagnostic and heuristic analysis based on the non-commutativity between diffusion and attention operators. This layer depends on a local linearization of attention and does not claim strict guarantees. Instead, it offers principled architectural guidance by quantifying geometric conflict between local smoothing and global routing.
This layered structure explicitly separates what is rigorously provable from what requires additional assumptions, thereby avoiding overstatement of theoretical claims.
3.2 Core Operator Definitions
We consider a sequence of length $L$ with hidden dimension $d$. Let $X\in\mathbb{R}^{L\times d}$ denote a layer representation, where $X_{i}\in\mathbb{R}^{d}$ is the feature vector at position $i$.
Boundary conditions.
To preserve information at sequence boundaries, we impose Neumann (zero-flux) boundary conditions via endpoint replication,
$x_{0}=x_{1},\qquad x_{L+1}=x_{L}.$
Discrete Neumann Laplacian.
For a vector $x\in\mathbb{R}^{L}$, the discrete Neumann Laplacian $\Delta_{N}$ acts elementwise as
$(\Delta_{N}x)_{i}=\begin{cases}x_{2}-x_{1},&i=1,\\ x_{i-1}-2x_{i}+x_{i+1},&2\leq i\leq L-1,\\ x_{L-1}-x_{L},&i=L.\end{cases}$ (1)
For matrix inputs, the operator is applied channel-wise along the sequence dimension, i.e., $(\Delta_{N}X)_{\cdot,j}=\Delta_{N}(X_{\cdot,j})$.
Single-scale diffusion.
The minimal diffusion module is defined as
$\mathcal{S}_{\alpha}(X)=(I+\alpha\Delta_{N})X,\qquad 0\leq\alpha<\tfrac{1}{2}.$
The constraint $\alpha<\tfrac{1}{2}$ acts as a stability budget, analogous to the CFL condition in explicit diffusion schemes, and guarantees spectral non-expansiveness and energy dissipation.
Multi-scale diffusion.
To accommodate heterogeneous feature lengths, we extend the operator to a multi-scale formulation by combining $K$ strided Laplacians with step sizes $h_{k}$:
$\mathcal{S}_{\boldsymbol{\alpha}}(X)=X+\sum_{k=1}^{K}\alpha_{k}\Delta_{N,h_{k}}X.$
A global budget constraint,
$\sum_{k=1}^{K}\alpha_{k}<\tfrac{1}{2},$
ensures that the same stability guarantees are preserved.
3.3 Systematic Insertion Design Space
We define a systematic design space consisting of seven canonical insertion strategies, ranging from early-stage pre-regularization to late-stage post-routing smoothing. These positions are summarized in Table 1.
Table 1: Canonical insertion strategies for PDE diffusion layers, spanning the full two-column width.
ID Strategy Composition Theoretical Expectation
P1 After-Embedding $X\leftarrow\mathcal{S}(\mathrm{Embed}(\cdot))$ Optimal: semantic manifold pre-regularization
P2 After-MLP $X\leftarrow\mathcal{S}(X+\mathrm{MLP}(X))$ Moderate: partial regularization
P3 Layer-Diffusion $X_{\ell+1}\leftarrow\mathcal{S}(\mathrm{Block}(X_{\ell}))$ Moderate: inter-layer smoothing
P4 Before-LN $X\leftarrow\mathrm{LN}(\mathcal{S}(X))$ Moderate: interaction with normalization
P5 In-Attention $\mathrm{Attn}(Q,K,V)\!\rightarrow\!\mathrm{Attn}(Q,K,\mathcal{S}(V))$ Unclear: internal intervention
P6 Head-Diffusion $H\leftarrow\mathcal{S}_{\text{head}}(H)$ Unclear: cross-head exchange
P7 After-Attention $X\leftarrow\mathcal{S}(X+\mathrm{Attn}(X))$ Worst: double-mixing conflict
The central contrast lies between early and late insertion. Strategy P1 smooths representations before global routing, allowing attention to aggregate information on a less noisy semantic manifold. In contrast, P7 applies local smoothing after non-local mixing, potentially blurring structures that attention has just constructed.
3.4 Implementation Details
The PDE diffusion operator has linear complexity in both time and memory. A single-scale diffusion requires $\mathcal{O}(Ld)$ operations, while the multi-scale variant scales as $\mathcal{O}(KLd)$.
To enforce the global stability budget exactly during training without manual clipping, we reparameterize the diffusion weights as
$\alpha_{k}=\tfrac{1}{2}\,\sigma(\eta)\,\mathrm{softmax}(\boldsymbol{\theta})_{k},$
where $\eta\in\mathbb{R}$ and $\boldsymbol{\theta}\in\mathbb{R}^{K}$ are unconstrained learnable parameters.
Depending on architectural choice, diffusion can be applied either directly or followed by normalization,
$\text{Out}=\mathcal{S}(X),\qquad\text{or}\qquad\text{Out}=\mathrm{LN}(\mathcal{S}(X)).$
3.5 Analysis Quantities: Roughness and Conflict Diagnostics
Local roughness.
We quantify local geometric irregularity along the sequence dimension using the discrete Dirichlet energy,
$\mathcal{E}(X)=\tfrac{1}{2}\langle X,-\Delta_{N}X\rangle=\tfrac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}\bigl(X_{i+1,j}-X_{i,j}\bigr)^{2}.$
This quantity measures how rapidly representations vary between neighboring tokens.
Geometric conflict.
To diagnose incompatibility between diffusion and attention, we define the diffusion–attention commutator,
$\mathcal{S},\mathcal{A}](X)=\mathcal{S}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}(X)).$
This operator measures the representational discrepancy between pre-attention and post-attention diffusion. In this work, we use it as a conceptual diagnostic for architectural compatibility (Layer III), rather than as a measured training-time signal.
4 Operator-Theoretic Analysis
This section develops an operator-theoretic account of how locality-inducing PDE diffusion interacts with self-attention. To avoid overclaiming, we separate results into three layers: (i) unconditional properties of the PDE diffusion operator in isolation; (ii) compositional perturbation bounds under standard local regularity assumptions for Transformer modules; and (iii) a diagnostic interpretation of diffusion–attention non-commutativity based on a local linearization of attention. Full proofs are deferred to Appendix B.
4.1 Layer I: Intrinsic Stability of Diffusion
We begin with properties that hold independently of the Transformer backbone. Let $\Delta_{N}\in\mathbb{R}^{L\times L}$ denote the discrete Neumann Laplacian (Eq. (1)), and define the single-step diffusion operator
$\mathcal{S}_{\alpha}\triangleq I+\alpha\Delta_{N},\qquad 0\leq\alpha<\tfrac{1}{2},$ (2)
applied channel-wise to $X\in\mathbb{R}^{L\times d}$.
Lemma 4.1 (Spectrum of the Discrete Neumann Laplacian).
The matrix $\Delta_{N}$ is real symmetric and negative semi-definite. Its eigenvalues are
$\lambda_{k}=-4\sin^{2}\!\left(\frac{\pi k}{2L}\right),\quad k=0,1,\ldots,L-1,$ (3)
where $\lambda_{0}=0$ (constant mode) and $\lambda_{k}\in[-4,0]$ for all $k$.
Theorem 4.2 (Spectral Non-Expansiveness and Strict Contraction).
For any $X\in\mathbb{R}^{L\times d}$ and $0\leq\alpha<\tfrac{1}{2}$,
$\lVert\mathcal{S}_{\alpha}\rVert_{2}\leq 1,\qquad\lVert\mathcal{S}_{\alpha}(X)\rVert_{F}\leq\lVert X\rVert_{F}.$ (4)
Let $\Pi_{\perp}$ be the orthogonal projection onto $\mathrm{Null}(\Delta_{N})^{\perp}$ (equivalently, the zero-mean subspace along the sequence dimension). Then
$\lVert\Pi_{\perp}\mathcal{S}_{\alpha}(X)\rVert_{F}$ $\leq\kappa(\alpha,L)\,\lVert\Pi_{\perp}X\rVert_{F},$ (5)
$\kappa(\alpha,L)$ $=\max_{k\geq 1}\bigl(1+\alpha\lambda_{k}\bigr)<1.$
The constraint $\alpha<\tfrac{1}{2}$ is a discrete stability budget analogous to the CFL condition for explicit diffusion: it ensures non-expansiveness and strict attenuation of non-constant (high-frequency) components, while leaving the constant mode unchanged.
To quantify locality, we use the discrete Dirichlet energy.
Definition 4.3 (Discrete Dirichlet Energy).
For $X\in\mathbb{R}^{L\times d}$,
$\mathcal{E}(X)\triangleq\frac{1}{2}\langle X,-\Delta_{N}X\rangle=\frac{1}{2}\sum_{j=1}^{d}\sum_{i=1}^{L-1}(X_{i+1,j}-X_{i,j})^{2},$ (6)
which measures local roughness along the sequence axis.
Theorem 4.4 (Monotone Energy Dissipation).
For any $0\leq\alpha<\tfrac{1}{2}$ and $X\in\mathbb{R}^{L\times d}$,
$\mathcal{E}(\mathcal{S}_{\alpha}(X))\leq\mathcal{E}(X),$ (7)
with strict inequality whenever $\Pi_{\perp}X\neq 0$ and $\alpha>0$.
Corollary 4.5 (Multi-Scale Stability Under a Global Budget).
Let $\mathcal{S}_{\boldsymbol{\alpha}}$ be the multi-scale diffusion operator defined above. If $\sum_{k=1}^{K}\alpha_{k}<\tfrac{1}{2}$, then $\mathcal{S}_{\boldsymbol{\alpha}}$ is spectrally non-expansive and dissipates $\mathcal{E}(\cdot)$ monotonically.
Practical takeaway.
With a strict global budget ($\alpha<\tfrac{1}{2}$ or $\sum_{k}\alpha_{k}<\tfrac{1}{2}$), PDE diffusion is a safe architectural primitive: it cannot amplify representations and it monotonically reduces local roughness.
4.2 Layer II: Compositional Perturbation Bounds
We next study how diffusion perturbs a Transformer computation when inserted at different locations. Let $f_{\ell}:\mathbb{R}^{L\times d}\to\mathbb{R}^{L\times d}$ denote sub-modules (LayerNorm, attention, MLP), and consider the composition $F=f_{m}\circ\cdots\circ f_{1}$.
Assumption 4.6 (Local Lipschitz Regularity).
For each $\ell$, there exists $L_{\ell}>0$ such that in a neighborhood of interest,
$\lVert f_{\ell}(X)-f_{\ell}(Y)\rVert_{F}\leq L_{\ell}\lVert X-Y\rVert_{F}.$ (8)
Define the variant with diffusion inserted after $f_{j}$:
$F^{(j)}\triangleq f_{m}\circ\cdots\circ f_{j+1}\circ\mathcal{S}_{\alpha}\circ f_{j}\circ\cdots\circ f_{1}.$ (9)
Theorem 4.7 (Perturbation Propagation Bound).
Under Assumption 4.6, for any input $X$ and $Z_{j}=(f_{j}\circ\cdots\circ f_{1})(X)$,
$\lVert F^{(j)}(X)-F(X)\rVert_{F}$ $\leq\left(\prod_{\ell=j+1}^{m}L_{\ell}\right)\alpha\,\lVert\Delta_{N}Z_{j}\rVert_{F}$ (10)
$\leq\left(\prod_{\ell=j+1}^{m}L_{\ell}\right)\alpha\,\sqrt{2\mathcal{E}(Z_{j})}.$
This bound isolates two insertion-dependent quantities: the local roughness at the insertion point, captured by $\mathcal{E}(Z_{j})$, and the downstream sensitivity, captured by $\prod_{\ell=j+1}^{m}L_{\ell}$. In particular, diffusion is guaranteed to be a small perturbation whenever both $\mathcal{E}(Z_{j})$ and the downstream amplification are controlled; conversely, sensitivity to insertion position is expected when either factor is large.
Practical takeaway.
The same diffusion operator can help or hurt depending on where it is inserted: its effect scales with insertion-point roughness and downstream amplification, explaining why early insertion is typically most beneficial.
4.3 Layer III: Diffusion–Attention Non-Commutativity as a Diagnostic
Finally, we formalize the difference between pre-attention and post-attention diffusion through an exact commutator identity. Let $\mathcal{A}$ denote the (possibly input-dependent) self-attention operator.
Proposition 4.8 (Commutator Identity).
For any $X$, define
$Y_{\mathrm{pre}}$ $\triangleq\mathcal{A}(\mathcal{S}_{\alpha}(X)),$ (11)
$Y_{\mathrm{post}}$ $\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X)).$
Then
$Y_{\mathrm{post}}-Y_{\mathrm{pre}}$ $=[\mathcal{S}_{\alpha},\mathcal{A}]X$ (12)
$\triangleq\mathcal{S}_{\alpha}(\mathcal{A}(X))-\mathcal{A}(\mathcal{S}_{\alpha}(X)),$
and hence $\lVert Y_{\mathrm{post}}-Y_{\mathrm{pre}}\rVert_{F}=\lVert[\mathcal{S}_{\alpha},\mathcal{A}]X\rVert_{F}$.
To interpret this quantity, we use a local linearization at a fixed input $X$: approximate attention as a mixing operator along the sequence dimension, $\mathcal{A}(X)\approx A(X)X$, where $A(X)\in\mathbb{R}^{L\times L}$ is row-stochastic. Substituting into the commutator yields the heuristic form
$\mathcal{S}_{\alpha},\mathcal{A}]X\approx\alpha\bigl(\Delta_{N}A(X)-A(X)\Delta_{N}\bigr)X,$ (13)
which makes explicit that non-commutativity arises from a mismatch between local geometry ($\Delta_{N}$) and non-local routing ($A(X)$). We emphasize that Eq. (13) is a diagnostic approximation rather than a global guarantee: self-attention is nonlinear and input-dependent, so this layer provides interpretable guidance, not a universal bound.
Practical takeaway.
Large diffusion–attention mismatch (a large commutator) indicates structural conflict between local smoothing and non-local routing, and motivates avoiding post-attention diffusion when attention induces strong non-local mixing.
4.4 Design Implications
The analysis suggests three practical implications. First, PDE diffusion provides a certified local geometric prior (Layer I) that is stable under a strict budget $\alpha<\tfrac{1}{2}$. Second, the effect of inserting PDE diffusion is controlled by insertion-point roughness and downstream amplification (Layer II), giving a principled notion of “position sensitivity” without relying on empirical observations. Third, the commutator $\mathcal{S}_{\alpha},\mathcal{A}]$ exactly measures the representational discrepancy between pre- and post-attention diffusion, and its linearized form attributes large discrepancies to incompatible local and non-local mixing patterns (Layer III). Together, these statements justify preferring PDE diffusion placements that regularize representations before aggressive non-local routing, while treating post-routing smoothing as potentially disruptive when the induced commutator is large.